OpenStudy (anonymous):

How do arrange to get from 1/2(1+5√)−1/2(1−5√) to √5 what steps do I need to take to arrive at the answer.

OpenStudy (anonymous):

\[1/2 (1 + \sqrt{5}) - 1/2(1 - \sqrt{5})\]

OpenStudy (anonymous):

to the following

OpenStudy (anonymous):

= \[\sqrt{5}\]

OpenStudy (anonymous):

Can you please show me the steps to get there?

OpenStudy (asnaseer):

factor out the 1/2 from both bracketed expressions

OpenStudy (anonymous):

Thanks. And the minus sign changes the second part?

OpenStudy (asnaseer):

yes - a similar expression might be:\[a(1+b)-a(1-b)=a[(1+b)-(1-b)]=a[1+b-1+b]=a[2b]=2ab\]

OpenStudy (anonymous):

thanks that is really helpful.

OpenStudy (asnaseer):

yw :)